write the slope-intercept equation of the line that is parallel to -3x=y+6 and has the same y-intersept as the graph -7x+4y=12
Answer by mananth(16946)

write the slope-intercept equation of the line that is parallel to -3x=y+6 and has the same y-intersept as the graph -7x+4y=12
-3x=y+6
y=-3x-6
slope = -3
..
A line parallel to this line will have same slope = -3
7x+4y=12
4y=-7x+12
/4
y=-7/4 +3
y intercept = 3
slope = -3, y intercept = 3
equation is
y = -3x+3